OpenStudy (nthenic_oftime):

Use even and odd properties of the trigonometric functions to find the exact value of the expression. csc(-Pi/6)

OpenStudy (anonymous):

\[\csc \left( -\frac{ \pi }{ 6 } \right)=-\csc \frac{ \pi }{ 6 }=\frac{ -1 }{ \sin \frac{ \pi }{ 6 } }=\frac{ -1 }{ \frac{ 1 }{ 2 } }=-2\]

OpenStudy (nthenic_oftime):

thank you :)